OpenStudy (anonymous):

What is the solution to the system of linear equations 3x+4y=18 x+3y=11 @Tawenger

OpenStudy (anonymous):

@kyrea do you know how to do simultaneous equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

Subtract the equations to get: 2x +y = 7 Solve for y y = 7 - 2x

OpenStudy (anonymous):

plug y into another equation, let's plug it into eq 2. you get x + 3(7-2x) = 11

OpenStudy (anonymous):

there are nun

OpenStudy (anonymous):

solve for x, x= 2

OpenStudy (anonymous):

plug in x into our other equation to get y =7-4 = 3

OpenStudy (anonymous):

of u make x the subject in the second equation you can substitute the value of x in that one into the first equation to get y

OpenStudy (anonymous):

so we get, x=2 and y = 3

OpenStudy (anonymous):

so for x+3y=11 x=11-3y

OpenStudy (anonymous):

now use that in 3x+4y=18 so it will be 3(11-3y)+4y=18

OpenStudy (anonymous):

of u expand the bracket and collect the like terms you can now solve for y can you do that @kyrea

OpenStudy (anonymous):

sure but i got the answer